OpenStudy (anonymous):

Write an equation in standard form of the line that passes through the given point and has the given slope m (–4, 3); m = 1/2

OpenStudy (mathstudent55):

Since you have the slope, start with the slope-intercept form of a line: y = mx + b, where m = slope. Plug in m = 1/2, x = -4, and y = 3. Solve for b. Then put it in the form ax + by + c = 0

OpenStudy (anonymous):

@mathstudent55 so b=1.5?

OpenStudy (mathstudent55):

y = mx + b 3 = (1/2)(-4) + b 3 = -2 + b 5 = b b = 5

OpenStudy (anonymous):

ohh ok i see now thanks so much

OpenStudy (mathstudent55):

Now plug in b into y = (1/2)x + b Then move all terms to the left equaling zero. Write it in the form ax + by + c = 0

OpenStudy (anonymous):

@mathstudent55 so (1/2)x+y = 5?

OpenStudy (mathstudent55):

y = (1/2)x + 5 Subtract both (1/2)x and 5 to the left side: -(1/2)x + y - 5 = 0 Multiply both sides by -2: x - 2y + 10 = 0

OpenStudy (anonymous):

ok thanks